OpenStudy (anonymous):

1.)An initial investment of $10,000 grows at 11% per year.What function represents the value of the investment after t years? 2.)Evaluate log 0.01 by using mental math?

OpenStudy (wolf1728):

1) The value of an investment is TOTAL = principal*(1+ rate)^t where t = years TOTAL = 10,000 * (1.11)^years 2) I'm assuming it is a base 10 log. So, whenever a base 10 log is to be found, if the number is exactly a power of 10, (.1, 1 , 10 , 100, 1,000) then the log is the exponent of that number .1 = -1, 1 = 0 10= 1 100= 2 1,000 = 3 So, the base 10 log of .01 is -2, because 10^-2 = .01

OpenStudy (anonymous):

ok thankyou! Could you also help me with this problem please? \[\log_{6} x ^{8}-\log_{6} x ^{5}=31\]

OpenStudy (wolf1728):

Well at least I can give it a try.

OpenStudy (wolf1728):

(log 10) x^8 - (log 10) x^5 = y If x=10, then 8 -5 =3 then y = 3 Well that's with base 10 logs and knowing the value of x. That's about all I can do.

jimthompson5910 (jim_thompson5910):

\[\Large \log_{6}(x^8)-\log_{6}(x^5) = 31\] \[\Large 8\log_{6}(x)-5\log_{6}(x) = 31\] \[\Large (8-5)\log_{6}(x) = 31\] \[\Large 3\log_{6}(x) = 31\] \[\Large \log_{6}(x) = \frac{31}{3}\] \[\Large x = 6^{31/3}\]

OpenStudy (anonymous):

okay! Thanks !!!

OpenStudy (wolf1728):

Very good Jim!!

OpenStudy (anonymous):

Do you guys know any good example websites for Math ?

jimthompson5910 (jim_thompson5910):

here is one page that shows various log rules http://www.purplemath.com/modules/logrules.htm

OpenStudy (anonymous):

thankyou! (:

jimthompson5910 (jim_thompson5910):

np

OpenStudy (wolf1728):

Mhm120 My own site www.1728.org has lots of math examples and calculators. Not too much on logs but here is a great log calculator http://1728.org/logrithm.htm

OpenStudy (wolf1728):

Incidentally Jim I tested out your answer 109,874,333.58 (or 6^31/3) and the log 6 of that number to the 8th power - the log 6 of that number to the 5th power = 31

OpenStudy (anonymous):

@jim_thompson5910 Could you please what you did for the log problem in words please?